OpenStudy (zackattack350):

Please help im lost. If a number is added to the numerator of 11/61 and twice the number is added to the denominator of 11/61, the resulting fraction is equivalent to 1/5.Find the number.

OpenStudy (lochana):

good one!. first you can define a variable for the unknown term. that is the initial number in your case. say x is our number. now you can write an equation out of given information

OpenStudy (zackattack350):

so using X we can make the equation into... x11/61+1/X2=1/5?

OpenStudy (lochana):

\[\frac{ x+\frac{ 11 }{ 61 } }{ 2x + \frac{ 11 }{ 61} } = \frac{ 1 }{ 5 }\]

OpenStudy (lochana):

too confusing ? that's what you are looking for. now you must solve for x.

OpenStudy (zackattack350):

let me try and calculate it really fast

OpenStudy (lochana):

okay

OpenStudy (zackattack350):

-44/183?

OpenStudy (lochana):

yep. that's what i got too.

OpenStudy (zackattack350):

ok thank you!

OpenStudy (lochana):

bye